OpenStudy (anonymous):

if |z|>= 3 the least value of |z + 1/z| is?

OpenStudy (anonymous):

As, IzI>=3 Iz+1/zI=IzI + I1/zI >=3+1/3=10/3

OpenStudy (amistre64):

\[\lim_{z\to\ inf}\frac{z+1}{z}=1\]just a thought

OpenStudy (anonymous):

Is it 

OpenStudy (anonymous):